# Defending Against Prompt Injection with DataFilter

### v1

@@ VI Conclusion and Discussions
Our work shows that it is possible to defend a black-box commercial LLM and preserve its utility by using another trained LLM to filter malicious injections from the data. DataFilter delivers a good balance of security, utility, and deployability. Even though it is trained only on basic attacks, it generalizes effectively to more complex injection strategies. Across multiple benchmarks, DataFilter consistently reduces attack success rates to near zero, outperforming detection- and prompt-based defenses, which either over-refuse benign inputs or miss attacks. Unlike system-level defenses, it requires no redesign of the agent or application and can be deployed in a plug-and-play manner to both commercial and open-weight models. Most importantly, DataFilter achieves these gains without sacrificing utility, maintaining task performance within a few percentage points (2%) of the undefended model. Together, these findings confirm that DataFilter is the first model-agnostic defense to simultaneously satisfy all three desiderata outlined in Section˜IV.
@@
Limitations. Our method still has below limitations. First, DataFilter introduces additional inference overhead, since the filter must run whenever new untrusted data is received. Second, we do not consider optimization-based adaptive attacks. They are more difficult to mount, particularly against proprietary models, where we expect our defense will be primarily used. We hypothesize that recent strong attacks [12, 18] can break our defense as it breaks all existing defenses if given sufficient computation. Third, while deployment is lightweight, some effort is still required from agent developers. In particular, DataFilter struggles with very long benign user prompts. Therefore, applications that use very long user messages should provide the filter message with a more concise user command, rather than the full user message. For example, in InjecAgent [22], the user message contains the user’s actual query together with tool introductions, example calls, and policies. Our filter model performs poorly if provided the entire user message but performs well if given the user’s query. Developers must therefore extract the short user instruction and pass it to DataFilter. Although this effort is modest, it does add an extra integration step compared to defenses fully embedded in the model.
@@ Acknowledgments
This work was supported by the KACST-UC Berkeley Center of Excellence for Secure Computing, the NSF ACTION center through NSF grant 2229876, and by generous gifts from Google, Meta, and Noyce foundation.

### v2

@@ VI Conclusion and Discussions
Our work shows that it is possible to defend a black-box commercial LLM and preserve its utility by using another trained LLM to filter malicious injections from the data. DataFilter delivers a good balance of security, utility, and deployability. Even though it is trained only on basic attacks, it generalizes effectively to more complex injection strategies. Similarly, our method transfers well to unseen domains: trained on Cleaned-Alpaca [47] (a single-turn instruction-tuning dataset), it generalizes to agentic benchmarks [66, 13] involving multi-turn tool calls in sandbox environments. Across multiple benchmarks, DataFilter consistently reduces attack success rates to near zero, outperforming detection- and prompt-based defenses, which either over-refuse benign inputs or miss attacks. Unlike system-level defenses, it requires no redesign of the agent or application and can be deployed in a plug-and-play manner to both commercial and open-weight models. Most importantly, DataFilter achieves these gains without sacrificing utility, maintaining task performance within a few percentage points (2%) of the undefended model. Together, these findings confirm that DataFilter is the first model-agnostic defense to simultaneously satisfy all three desiderata outlined in Section˜IV.
@@
Limitations. Our method still has below limitations. First, DataFilter introduces additional inference overhead, since the filter must run whenever new untrusted data is received. Second, our defense cannot defend against the strong optimization-based adaptive attacks. As discussed in Section VIII, a recent strong attack [33] breaks our defense, as it breaches all existing defenses. Third, while deployment is lightweight, some effort is still required from agent developers. In particular, DataFilter struggles with very long benign user prompts. Therefore, applications that use very long user messages should provide the filter message with a more concise user command, rather than the full user message. For example, in InjecAgent [66], the user message contains the user’s actual query together with tool introductions, example calls, and policies. Our filter model performs poorly if provided the entire user message but performs well if given the user’s query. Developers must therefore extract the short user instruction and pass it to DataFilter. Although this effort is modest, it does add an extra integration step compared to defenses fully embedded in the model.
@@ Acknowledgments
This work was supported by the KACST-UC Berkeley Center of Excellence for Secure Computing, the NSF ACTION center through NSF grant 2229876, and by generous gifts from Google, Meta, and Noyce foundation. We thank Chawin Sitawarin for providing the results of the adaptive attack reported in Table VIII.